# Confidence-based Estimators for Predictive Performance in Model Monitoring

### v1

@@ Abstract
Recently, novel methods for estimating the predictive performance of a model when ground truth is unavailable have been developed. Many of these methods leverage model confidence or other uncertainty estimates and are experimentally compared against a naive baseline method, namely Average Confidence (AC), which estimates model accuracy as the average of confidence scores for a given set of predictions. However, until now the theoretical properties of the AC method have not been properly explored. In this paper, we try to fill this gap by reviewing the AC method and show that under certain general assumptions, it is an unbiased and consistent estimator of model accuracy with many desirable properties. We also compare this baseline estimator against some more complex estimators empirically and show that in many cases the AC method is able to beat the others, although the comparative quality of the different estimators is heavily case-dependent.
@@ 1 Introduction
The last decade has seen a considerable improvement in the predictive capability of machine learning (ML) models. However, as techniques in designing and implementing these models have matured, their use in real-life scenarios is still in its infancy. There are many reasons for the slow adoption of ML models in industrial practices. The real-life use cases require the models to be integrated with other services and data pipelines, forming a complex Machine Learning system (Schröder and Schulz, 2022). Furthermore, questions of reliability need to be addressed.
@@
Monitoring can be directed towards many different stages in an ML system. Roughly speaking, one can monitor input data (Rabanser et al., 2019), the inner workings of a model (Myllyaho et al., 2022), or model outputs (Hendrycks and Gimpel, 2017; Corbière et al., 2019). The monitoring can be targeted to capture anomalies, drifts in the data, or performance deterioration (Klaise et al., 2020). One could even extend the scope of monitoring to capture aspects such as server throughput or even business impact. In this work, we focus only on methods that monitor the predictive performance of models based on their outputs.
@@
In some cases, an ML practitioner can compare the predictions of a model with the actual ground truth (GT) values with some acceptable latency. For example, if a model is used to predict tomorrow’s stock prices, one can assess the quality of those predictions with a lag of one day. In some cases, the GT values for a portion of predictions can be queried from an expert (Ginart et al., 2022). In such cases, monitoring predictive performance is relatively straightforward (Klaise et al., 2020) and one can choose from a wide range of algorithms (Bayram et al., 2022).
@@
Unfortunately, there are plenty of cases where GT becomes available only after a substantial lag or not at all. In such cases, one can not directly monitor the predictive performance of an ML system with acceptable latency. Historically in these situations, only shifts in the data have been monitored with the hope that they are indicative of shifts in model performance. However, this is not always the case, and making such assumptions might result in unnecessary re-trainings inducing unnecessary costs (Baier et al., 2021). This dilemma becomes even worse with high-dimensional data, although using techniques such as dimensionality reduction (Rabanser et al., 2019) can alleviate this to some extent. Clearly, tools for distinguishing meaningful shifts from redundant ones are needed.
@@
Only quite recently, there have been several propositions in how to solve the unsupervised accuracy estimation problem, where instead of monitoring shifts in the data, one tries to estimate the predictive performance directly (Baek et al., 2022; Chen et al., 2021a; Deng et al., 2023; Garg et al., 2022; Guillory et al., 2021; Jiang et al., 2022; Lu et al., 2023) in the absence of GT. Many of the papers proposing these new estimation methods use the Average Confidence (AC) (Hendrycks and Gimpel, 2017) of a model as a simple baseline for estimating the predictive accuracy and experimentally show their method to beat this baseline under different scenarios. However, none of these papers explain in detail, why they choose to use AC as a comparative baseline, nor present the theoretical properties of this baseline.
@@
In this paper, we present a justification for using AC as a baseline method for estimating the predictive performance of an ML system in detail. We prove theoretically that under certain assumptions, this method is an unbiased and consistent estimator of the model’s predictive performance. Furthermore, we complement and contrast some of the results presented in earlier papers (Guillory et al., 2021; Garg et al., 2022) by performing empirical tests under different scenarios.
@@ 2.1 Unsupervised accuracy estimation under dataset shift
Let us assume that we have trained a supervised ML model $f$ using data from a source distribution $p_{s}(\boldsymbol{x},y)$ available to us during development. Once the model is deployed, it will make predictions using data from a target distribution $p_{t}(\boldsymbol{x},y)$. Typically $p_{s}(\boldsymbol{x},y)\neq p_{t}(\boldsymbol{x},y)$. Following the terminology of Moreno-Torres et al. (2012), we label this inequality as dataset shift.
@@
For an estimator to work, assumptions about the nature of the shift have to be made (Garg et al., 2022). Since for any distribution $p(\boldsymbol{x},y)=p(y|\boldsymbol{x})p(\boldsymbol{x})$, the dataset shift can be caused by concept shift, where $p_{s}(y|\boldsymbol{x})\neq p_{t}(y|\boldsymbol{x})$. If this is not the case, it must be caused by covariate shift, where $p_{s}(\boldsymbol{x})\neq p_{t}(\boldsymbol{x})$. These causes are taken as mutually exclusive. If both occur, the generic term of dataset shift is used. Alternatively, one can also decompose $p(\boldsymbol{x},y)=p(\boldsymbol{x}|y)p(y)$. In this case, a situation where $p_{s}(\boldsymbol{x}|y)=p_{t}(\boldsymbol{x}|y)$ and $p_{s}(y)\neq p_{t}(y)$ is called label shift. In this paper, we focus in cases of covariate shift.
@@ 2.2 Traditional process monitoring
Statistical Process Control (SPC) was pioneered by Shewhart at Bell Labs in the 1920s for systematic monitoring and quality control of industrial production processes. This approach has since been applied to all kinds of processes outside manufacturing with great success. In this wider context, a process can be defined as "everything required to turn an input into an output for a customer" (Stapenhurst, 2005). As such, the predictions of an ML model fit this pattern as well.
@@
Control charts are central tools used to monitor processes in SPC. A control chart is essentially a run chart of some statistic calculated from consecutive samples. It tracks whether the variation of the statistic from the mean stays within preset control limits. There are various established sets of decision rules (such as the Western Electric rules, see e.g. (Montgomery, 2009)) for recognizing when the data shown on a control chart implies that the process has likely strayed outside statistical control. If this is the case, an alarm is raised for the user. In an ML monitoring setting, widely-used algorithms such as Drift Detection Method (DDM) and its numerous variants (Bayram et al., 2022) are implementations of SPC. However, these methods can only be applied with access to GT.
@@ 2.3 Confidence calibration
In addition to predicting a certain class, most classifier models also output (or can be modified to output) a vector of soft scores with one value for each class. Usually, the values of the vectors are normalized on the interval $[0,1]$, summing to one, which means they form a categorical probability distribution over the label space. The maximum value of the normalized vector can be interpreted as a confidence score (Guo et al., 2017) for the prediction, where the value $1$ signals maximal confidence.
In addition to classifiers, confidence scores can be derived for many other model types. Recently, one increasingly popular approach with non-classifier models has been to train an auxiliary model to assign a confidence score for the predictions of the base model (Chen et al., 2019; Corbière et al., 2019; DeVries and Taylor, 2018; Kivimäki et al., 2023; Shao et al., 2020). Using such techniques, one can attach a confidence score to practically any type of model as long as criteria for treating each prediction of the base model as either conforming or non-conforming exists.
In the context of classifier models, it is tempting to interpret a confidence score as the probability that a given instance was predicted correctly. However, it is rather common that the predicted probabilities do not align with empirical probabilities. That is, if one monitors all the predictions that were assigned a certain confidence score, the fraction of correct predictions among those predictions would deviate from the confidence score significantly (Guo et al., 2017; Zadrozny and Elkan, 2001). If the confidence scores produced by a model align with empirical probabilities, they are said to be calibrated. We will next provide a more formal definition of calibration.
@@
For the sake of brevity, in the rest of this paper, we will refer to the conforming predictions as correct and to the non-conforming as incorrect. In practice, no model can achieve perfect calibration, so for the concept of calibration to be meaningful, we must be able to quantify calibration error. How this should be done is an active discussion within the scientific community (see e.g. (Guo et al., 2017; Nixon et al., 2019; Posocco and Bonnefoy, 2021; Vaicenavicius et al., 2019)). One of the currently most used metrics is Expected Calibration Error (ECE), which requires the predictions to be binned into $M$ bins according to their confidence scores. The hyperparameter $M$ is set according to user preferences and affects the estimated calibration error. ECE is defined as
@@
where $n$ is the total number of predictions, $B_{m}$ the set of predictions in bin $m$, and $A_{m}$ and $C_{m}$ the average accuracy and confidence of predictions in bin $m$ respectively. In the standard version, the bins are assumed to be equiwidth. In this work, however, we choose to use equisize binning, where bin boundaries are adapted so that each bin holds the same number of predictions. This adaptive binning scheme gives rise to Adaptive Expected Calibration Error (ACE), which is shown to be more robust with respect to the number of bins hyperparameter (Nixon et al., 2019). We use $M=20$ in our experiments. Other than the choice of binning, ACE is calculated exactly the same way as ECE.
If the confidence scores are not calibrated, a post hoc calibration mapping can be applied to reduce the calibration error and make the confidence scores match empirical probabilities more accurately (e.g. (Alexandari et al., 2020; Guo et al., 2017; Kull et al., 2017, 2019; Kumar et al., 2019; Naeini et al., 2015; Platt, 1999; Zadrozny and Elkan, 2001, 2002)). There are also ways in training models inducing an implicit calibration effect on the confidence scores (e.g. (Kumar et al., 2018; Mukhoti et al., 2020; Müller et al., 2019; Ovadia et al., 2019; Seo et al., 2019; Zhang et al., 2020)), although most of these methods were not originally designed with calibration in mind.
@@ 3 Related Work
In recent years, there have been many propositions on how to perform unsupervised accuracy estimation. Some of these methods leverage model confidence expressed as the maximum of an output vector from the final softmax layer of a neural network (Guillory et al., 2021; Garg et al., 2022; Deng et al., 2023). Alternatively, the prediction entropy has been used as a measure of uncertainty. Others train several models and try to assess the predictive accuracy by tracking discrepancies in the predictions of those models (Baek et al., 2022; Chen et al., 2021a; Jiang et al., 2022). One approach is to use importance weighting (IW) (Lu et al., 2022) to account for the differences between the source and target distributions (Chen et al., 2021b). The methods presented by Guillory et al. (2021) and Garg et al. (2022) are most similar to AC, so we will explain them in more detail below and focus on them in our experiments.
@@
In the Difference of Confidences (DoC) method (Guillory et al., 2021), one first measures the difference of the average confidences of a model on a training set and on multiple shifted datasets, which are created from the training dataset by adding different kinds of synthetic noise. Then, an auxiliary regressor model is trained to estimate the base model’s accuracy on the shifted datasets using the difference in confidence scores between the training set and the shifted set as the only feature. The authors show experimentally that this approach leads to better generalization on both synthetic and natural shifts when compared to using common distributional distances such as Frechet distance (Fréchet, 1957) and Maximum Mean Discrepancy (Muandet et al., 2017) as features (Guillory et al., 2021). They also experiment with using the difference of average entropies (DoE) in place of average confidences arriving at similar results. Furthermore, and most interestingly for our current pursuit, they compare DoC against three simple regression-free baselines: Average Confidence (AC), Average Confidence after calibration with Temperature Scaling (AC TempScaling) (Guo et al., 2017), and subtracting the Difference of Confidences feature value from training accuracy (DoC-Feat). They show that AC is able to beat all of the regression models trained using distributional distances but loses the comparison against DoC, DoE, and Doc-Feat. Most interestingly, they claim that AC performs better than AC TempScaling, implying that calibration might be harmful.
@@
One shortcoming of the DoC approach is that it requires the creation of several shifted datasets. This might be infeasible in many cases. Furthermore, no guarantees can be given that the regression function learned with the synthetic shifted datasets would work in all cases of real-world shifts. In fact,  Garg et al. (2022) experimentally show the existence of cases where the DoC approach would fail. Instead of using the difference of confidence, they propose to use Average Thresholded Confidence (ATC), where one first finds a confidence threshold such that the fraction of training samples below that threshold equals the prediction error rate on the training set. Then, for a shifted dataset the estimated accuracy is the fraction of samples above the threshold. This method is shown to beat AC, DoC-Feat, IW and the GDE method described by Jiang et al. (2022). No direct comparison between ATC and DoC is made, with the rationale that ATC is regression-free and DoC is not. Contrary to Guillory et al. (2021), the authors found that calibration enhances the effectiveness of their method. In fact, for all methods they tested against (including AC and DoC-Feat), calibration improved the estimates in almost all of the test cases.
@@
On a more general note, all of the methods mentioned in the first paragraph of this section are tested on the dataset level. That is, a model is first trained on one dataset and its accuracy is evaluated on a shifted dataset, where the shifted dataset consists of thousands or even tens of thousands of samples. For the estimates to be useful in most practical monitoring settings, they should be relatively accurate for much smaller sets of samples, in the order of hundreds of samples. None of the methods are tested on such smaller sets of samples. To make matters worse, the variance of estimation error is not explored for any of the methods. Thus, even though the methods might give good estimates on a dataset level, the estimates on smaller batches of samples might express large variance, which could seriously hamper their usability in estimating accuracy in a monitoring setting. Furthermore, no theoretical guarantees for their methods are given by Baek et al. (2022); Deng et al. (2023); Guillory et al. (2021), and even where such guarantees are given, they depend on some highly specific assumptions. This raises the question of how well would the experimental results described in these papers hold in other settings.
@@
The methods are mainly tested in a multiclass image classification setting (although (Garg et al., 2022; Chen et al., 2021a; Lu et al., 2023) also include some tests with textual data) with neural networks and the confidence scores are taken to be the maximum of the softmax output. This typically results in uncalibrated confidence scores (Guo et al., 2017) and although all of the methods utilizing these confidence scores also report results for calibrated confidence scores, the calibration procedure is equated with temperature scaling (Guo et al., 2017) and the calibration errors for neither the original nor shifted datasets are reported. This makes using AC as a comparative baseline somewhat problematic since the performance of AC might be seriously hampered under miscalibration (as we shall soon see).
@@
Although it is taken as common knowledge that calibration error tends to increase under covariate shift (Ovadia et al., 2019), this is only shown to hold for softmax classifiers calibrated with temperature scaling (Guo et al., 2017). In fact, methods to ensure low calibration error under covariate shift are being developed (Białek et al., 2024). Interestingly, AC baseline is shown by (Guillory et al., 2021) to be comparable or better to DoC under training schemes, which contain data augmentations, such as AugMix (Hendrycks* et al., 2020) and DeepAugment (Hendrycks et al., 2021). Such augmentation schemes are known to produce implicit calibration for the trained models (Hendrycks* et al., 2020). This raises the question of whether the AC baseline would have been more performant under some other calibration scheme.

### v2

@@ Abstract
Recently, novel methods for estimating the predictive performance of a model when ground truth is unavailable have been developed. Many of these methods leverage model confidence or other uncertainty estimates and are experimentally compared against a naive baseline method, namely Average Confidence (AC), which estimates model accuracy as the average of confidence scores for a given set of predictions. However, until now the theoretical properties of the AC method have not been properly explored. In this paper, we bridge this gap by reviewing the AC method and show that under certain general assumptions, it is an unbiased and consistent estimator of model accuracy. We also augment the AC method by deriving valid confidence intervals for the estimates it produces. These contributions elevate AC from an ad-hoc estimator to a principled one, encouraging its use in practice.
We complement our theoretical results with empirical experiments, comparing AC against more complex estimators in a monitoring setting under covariate shift. We conduct our experiments using synthetic datasets, which allow for full control over the nature of the shift. Our experiments with binary classifiers show that the AC method is able to beat other estimators in many cases. However, the comparative quality of the different estimators is found to be heavily case-dependent.
@@ 1 Introduction
The last decade has seen a considerable improvement in the predictive capability of machine learning (ML) models. However, as techniques in designing and implementing these models have matured, their use in real-life scenarios is still in its infancy. There are many reasons for the slow adoption of ML models in industrial practices. The real-life use cases require the models to be integrated with other services and data pipelines, forming a complex Machine Learning system (?). Furthermore, questions of reliability need to be addressed.
@@
Monitoring can be directed towards many different stages in an ML system. Roughly speaking, one can monitor input data (?), the inner workings of a model (?), or model outputs (?, ?). The monitoring can be targeted to capture anomalies, drifts in the data, or performance deterioration (?). One could even extend the scope of monitoring to capture aspects such as server throughput or even business impact. In this work, we focus only on methods that monitor the predictive performance of models based on their outputs.
@@
In some cases, an ML practitioner can compare the predictions of a model with the actual ground truth (GT) values with some acceptable latency. For example, if a model is used to predict tomorrow’s stock prices, one can assess the quality of those predictions with a lag of one day. In some cases, the GT values for a portion of predictions can be queried from an expert (?). In such cases, monitoring predictive performance is relatively straightforward (?) and one can choose from a wide range of algorithms (?).
@@
Unfortunately, there are plenty of cases where GT becomes available only after a substantial lag or not at all. In such cases, one can not directly monitor the predictive performance of an ML system with acceptable latency. Historically in these situations, only shifts in the data have been monitored with the hope that they are indicative of shifts in model performance. However, this is not always the case, and making such assumptions might result in unnecessary re-trainings inducing unnecessary costs (?). This dilemma becomes even worse with high-dimensional data, although using techniques such as dimensionality reduction (?) can alleviate this to some extent. Clearly, tools for distinguishing meaningful shifts from redundant ones are needed.
@@
Only quite recently, there have been several propositions in how to solve the unsupervised accuracy estimation problem, where instead of monitoring shifts in the data, one tries to estimate the predictive performance directly (?, ?, ?, ?, ?, ?, ?) in the absence of GT. Many of the papers proposing these new estimation methods use the Average Confidence (AC) (?) of a model as a simple baseline for estimating the predictive accuracy and experimentally show their method to beat this baseline under different scenarios. However, none of these papers explain in detail, why they choose to use AC as a comparative baseline, nor present the theoretical properties of this baseline.
@@
In this paper, we present a justification for using AC as a baseline method for estimating the predictive performance of an ML system in detail. We prove theoretically that under certain assumptions, this method is an unbiased and consistent estimator of the model’s predictive performance. Furthermore, we complement and contrast some of the results presented in earlier papers (?, ?) by performing empirical tests under different scenarios.
@@ 2.1 Unsupervised Accuracy Estimation under Dataset Shift
Let us assume that we have trained a supervised ML model $f$ using data from a source distribution $p_{s}(\boldsymbol{x},y)$ available to us during development. Once the model is deployed, it will make predictions using data from a target distribution $p_{t}(\boldsymbol{x},y)$. Typically $p_{s}(\boldsymbol{x},y)\neq p_{t}(\boldsymbol{x},y)$. Following the terminology of ? (?), we label this inequality as dataset shift.
@@
For an estimator to work, assumptions about the nature of the shift have to be made (?). Since for any distribution $p(\boldsymbol{x},y)=p(y|\boldsymbol{x})p(\boldsymbol{x})$, the dataset shift can be caused by concept shift, where $p_{s}(y|\boldsymbol{x})\neq p_{t}(y|\boldsymbol{x})$. If this is not the case, it must be caused by covariate shift, where $p_{s}(\boldsymbol{x})\neq p_{t}(\boldsymbol{x})$. These causes are taken as mutually exclusive. If both occur, the generic term of dataset shift is used. Alternatively, one can also decompose $p(\boldsymbol{x},y)=p(\boldsymbol{x}|y)p(y)$. In this case, a situation where $p_{s}(\boldsymbol{x}|y)=p_{t}(\boldsymbol{x}|y)$ and $p_{s}(y)\neq p_{t}(y)$ is called label shift. In this paper, we focus on cases of covariate shift.
@@ 2.2 Traditional Process Monitoring
Statistical Process Control (SPC) was pioneered by Shewhart at Bell Labs in the 1920s for systematic monitoring and quality control of industrial production processes. This approach has since been applied to all kinds of processes outside manufacturing with great success. In this wider context, a process can be defined as ”everything required to turn an input into an output for a customer” (?). As such, the predictions of an ML model fit this pattern as well.
@@
Control charts are central tools used to monitor processes in SPC. A control chart is essentially a run chart of some statistic calculated from consecutive samples. It tracks whether the variation of the statistic from the mean stays within preset control limits. There are various established sets of decision rules (such as the Western Electric rules, see e.g. ?) for recognizing when the data shown on a control chart implies that the process has likely strayed outside statistical control. If this is the case, an alarm is raised for the user. In an ML monitoring setting, widely-used algorithms such as Drift Detection Method (DDM) and its numerous variants (?) are implementations of SPC. However, these methods can only be applied with access to GT.
@@ 2.3 Confidence Calibration
In addition to predicting a certain class, most classifier models also output (or can be modified to output) a vector of soft scores with one value for each class. Usually, the values of the vectors are normalized on the interval $[0,1]$, summing to one, which means they form a categorical probability distribution over the label space. The maximum value of the normalized vector can be interpreted as a confidence score (?) for the prediction, where the value $1$ signals maximal confidence.
In addition to classifiers, confidence scores can be derived for many other model types. Recently, one increasingly popular approach with non-classifier models has been to train an auxiliary model to assign a confidence score for the predictions of the base model (?, ?, ?, ?, ?). Using such techniques, one can attach a confidence score to practically any type of model as long as criteria for treating each prediction of the base model as either conforming or non-conforming exists.
In the context of classifier models, it is tempting to interpret a confidence score as the probability that a given instance was predicted correctly. However, it is rather common that the predicted probabilities do not align with empirical probabilities. That is, if one monitors all the predictions that were assigned a certain confidence score, the fraction of correct predictions among those predictions would deviate from the confidence score significantly (?, ?). If the confidence scores produced by a model align with empirical probabilities, they are said to be calibrated. We will next provide a more formal definition of calibration.
@@
For the sake of brevity, in the rest of this paper, we will refer to the conforming predictions as correct and to the non-conforming as incorrect. In practice, no model can achieve perfect calibration, so for the concept of calibration to be meaningful, we must be able to quantify calibration error. How this should be done is an active discussion within the scientific community  (see e.g., ?, ?, ?, ?). One of the currently most used metrics is Expected Calibration Error (ECE), which requires the predictions to be binned into $M$ bins according to their confidence scores. The hyperparameter $M$ is set according to user preferences and affects the estimated calibration error. ECE is defined as
@@
where $n$ is the total number of predictions, $B_{m}$ the set of predictions in bin $m$, and $A_{m}$ and $C_{m}$ the average accuracy and confidence of predictions in bin $m$ respectively. In the standard version, the bins are assumed to be equiwidth. In this work, however, we choose to use equisize binning, where bin boundaries are adapted so that each bin holds the same number of predictions. This adaptive binning scheme gives rise to Adaptive Expected Calibration Error (ACE), which is shown to be more robust with respect to the number of bins hyperparameter (?). We use $M=20$ in our experiments. Other than the choice of binning, ACE is calculated exactly the same way as ECE.
If the confidence scores are not calibrated, a post hoc calibration mapping can be applied to reduce the calibration error and make the confidence scores match empirical probabilities more accurately (e.g., ?, ?, ?, ?, ?, ?, ?, ?, ?). There are also ways in training models inducing an implicit calibration effect on the confidence scores (e.g., ?, ?, ?, ?, ?, ?), although most of these methods were not originally designed with calibration in mind.
@@ 3 Related Work
In recent years, there have been many propositions on how to perform unsupervised accuracy estimation. Some of these methods leverage model confidence expressed as the maximum of an output vector from the final softmax layer of a neural network (?, ?, ?). Alternatively, the prediction entropy has been used as a measure of uncertainty. Others train several models and try to assess the predictive accuracy by tracking discrepancies in the predictions of those models (?, ?, ?). One approach is to use importance weighting (IW) (?) to account for the differences between the source and target distributions (?). The methods presented by ? (?) and ? (?) are most similar to AC, so we will explain them in more detail below and focus on them in our experiments.
@@
In the Difference of Confidences (DoC) method (?), one first measures the difference of the average confidences of a model on a training set and on multiple shifted datasets, which are created from the training dataset by adding different kinds of synthetic noise. Then, an auxiliary regressor model is trained to estimate the base model’s accuracy on the shifted datasets using the difference in confidence scores between the training set and the shifted set as the only feature. The authors show experimentally that this approach leads to better generalization on both synthetic and natural shifts when compared to using common distributional distances such as Frechet distance (?) and Maximum Mean Discrepancy (?) as features (?). They also experiment with using the difference of average entropies (DoE) in place of average confidences arriving at similar results. Furthermore, and most interestingly for our current pursuit, they compare DoC against three simple regression-free baselines: Average Confidence (AC), Average Confidence after calibration with Temperature Scaling (AC TempScaling) (?), and subtracting the Difference of Confidences feature value from training accuracy (DoC-Feat). They show that AC is able to beat all of the regression models trained using distributional distances but loses the comparison against DoC, DoE, and Doc-Feat. Most interestingly, they claim that AC performs better than AC TempScaling, implying that calibration might be harmful.
@@
One shortcoming of the DoC approach is that it requires the creation of several shifted datasets. This might be infeasible in many cases. Furthermore, no guarantees can be given that the regression function learned with the synthetic shifted datasets would work in all cases of real-world shifts. In fact,  ? (?) experimentally show the existence of cases where the DoC approach would fail. Instead of using the difference of confidence, they propose to use Average Thresholded Confidence (ATC), where one first finds a confidence threshold such that the fraction of training samples below that threshold equals the prediction error rate on the training set. Then, for a shifted dataset the estimated accuracy is the fraction of samples above the threshold. This method is shown to beat AC, DoC-Feat, IW and the GDE method described by ? (?). No direct comparison between ATC and DoC is made, with the rationale that ATC is regression-free and DoC is not. Contrary to ? (?), the authors found that calibration enhances the effectiveness of their method. In fact, for all methods they tested against (including AC and DoC-Feat), calibration improved the estimates in almost all of the test cases.
@@
On a more general note, all of the methods mentioned in the first paragraph of this section are tested on the dataset level. That is, a model is first trained on one dataset and its accuracy is evaluated on a shifted dataset, where the shifted dataset consists of thousands or even tens of thousands of samples. For the estimates to be useful in most practical monitoring settings, they should be relatively accurate for much smaller sets of samples, in the order of hundreds of samples. None of the methods are tested on such smaller sets of samples. To make matters worse, the variance of estimation error is not explored for any of the methods. Thus, even though the methods might give good estimates on a dataset level, the estimates on smaller batches of samples might express large variance, which could seriously hamper their usability in estimating accuracy in a monitoring setting. Furthermore, no theoretical guarantees for their methods are given by ? (?, ?, ?), and even where such guarantees are given, they depend on some highly specific assumptions. This raises the question of how well would the experimental results described in these papers hold in other settings.
@@
The methods are mainly tested in a multiclass image classification setting (although ?, ?, ?, also include some tests with textual data) with neural networks, and the confidence scores are taken to be the maximum of the softmax output. This typically results in uncalibrated confidence scores (?) and although all of the methods utilizing these confidence scores also report results for calibrated confidence scores, the calibration procedure is equated with temperature scaling (?) and the calibration errors for neither the original nor shifted datasets are reported. This makes using AC as a comparative baseline somewhat problematic since the performance of AC might be seriously hampered under miscalibration (as we shall soon see).
@@
Although it is taken as common knowledge that calibration error tends to increase under covariate shift (?), this is only shown to hold for softmax classifiers calibrated with temperature scaling (?). In fact, methods to ensure low calibration error under covariate shift are being developed (?). Interestingly, AC baseline is shown by (?) to be comparable or better to DoC under training schemes, which contain data augmentations, such as AugMix (?) and DeepAugment (?). Such augmentation schemes are known to produce implicit calibration for the trained models (?). This raises the question of whether the AC baseline would have been more performant under some other calibration scheme.